Find the value of y so that the line passing through (9, 0) and (3, y) has a slope of -1/2
Answer by Maths68(1474)

Use Equation of line, Point-slope form which is
m=(y2-y1)/(x2-x1)
Given
(x1,y1)(x2,y2) = (9, 0)(3, y)
m = -1/2
Plug in the values in the equation, we have
m=(y2-y1)/(x2-x1)
-1/2=(y-0)/(3-9)
-1/2=y/-6
(-1/2)*(-6)=y
6/2=y
3=y
y=3